OpenStudy (anonymous):

In a 30ounces solution, 40% is alcohol. If 30 ounces of water is added to the solution, what percent of the new solution is alcohol?

OpenStudy (anonymous):

should be 20%

OpenStudy (anonymous):

No, no direct answers not needed. If you can explain the concept, only that would be helpful.

OpenStudy (anonymous):

40% of 30=12 Ounces alcohol 30 more ounces of water is added which means now there is 30+18=48 ounces of water percentage of alcohol = 12/60 =20%

OpenStudy (anonymous):

Thanks Hasham.

OpenStudy (anonymous):

Call me Hashir Khan :P and Yw.